OpenStudy (anonymous):

The equation y = 1.6x represents the number of laps Henry can swim over time, where y is the number of laps and x is time in minutes. The table shows the number of laps Larry can swim over time. How many laps can each boy complete per minute, and who can swim laps at a faster rate? Henry: laps per minute Larry: laps per minute The boy who swims at a faster rate is . (Type H for Henry or L for Larry) Larry Time (min) 2.5 4 5.5 8 Laps completed 4.5 7.2 9.9 14.4

OpenStudy (anonymous):

@jhonyy9

OpenStudy (aum):

For Henry: y = 1.6 * x The slope is 1.6 which means he can swim 1.6 laps per minute. For Larry: Time (min) 2.5 4 5.5 8 Laps completed 4.5 7.2 9.9 14.4 Differences in time: 1.5 1.5 2.5 Differences in laps: 2.7 2.7 4.5 Slope = 2.7/1.5 = 27/15 = 9/5 = 1.8 Larry does 1.8 laps per minute. Larry swims faster.

OpenStudy (anonymous):

@aum Thank you soooo much